OpenStudy (anonymous):

Kim rode her bike at a rate of 24km/h to the post office and 16km/h back home.  If her trip to the post office and back took a total of 75 minutes, how far does Kim live from the post office?

OpenStudy (valpey):

\[\frac{Distance}{24\frac{km}{h}} + \frac{Distance}{16\frac{km}{h}} = 75\ minutes\]

OpenStudy (valpey):

\[75\ minutes *\frac{1\ hour}{60\ minutes} = 1.25\ hours\]

OpenStudy (anonymous):

Then what happens

OpenStudy (valpey):

\[Distance*(\frac{1}{24}+\frac{1}{16}) = 1.25\](in kilometers)

OpenStudy (valpey):

Basically we have Distance divided by speed equals time. Time 1 + Time 2 = 1.25 hours and Distance 1 = Distance 2.

OpenStudy (anonymous):

Hi Hey good day

OpenStudy (valpey):

Can you add those fractions?

OpenStudy (anonymous):

It is 5/48

OpenStudy (valpey):

Good, so now multiply both sides of the equation by the reciprocal and you have the Distance to the post office in kilometers.

OpenStudy (valpey):

When you are done you should check to make sure that Time 1 and Time 2 make sense.

OpenStudy (anonymous):

How do you do that can u show me

OpenStudy (anonymous):

75 trimestre 5/48

OpenStudy (valpey):

\[Distance*\frac{5}{48}*\frac{48}{5} = 1.25*\frac{48}{5} \]

OpenStudy (anonymous):

Distàncies is75 rite

OpenStudy (valpey):

No. Distance is 1.25 * 48/5

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

So Ife You cursos divided You get 1.25 trimestre 48/5= that

OpenStudy (anonymous):

So 12 =12 is Thatcher t'he correct answer

OpenStudy (anonymous):

hello is that answer

OpenStudy (valpey):

\[Distance∗\Big(\frac{1}{24}+\frac{1}{12}\Big)=1.25\] \[Distance∗\frac{5}{48}=1.25\] \[Distance∗\frac{5}{48}∗\frac{48}{5}=1.25∗\frac{48}{5}\] \[Distance=1.25∗\frac{48}{5}=12\] Now check to make sure the times make sense. How long did it take to get to the post office?

OpenStudy (anonymous):

How did you get 1/12

OpenStudy (valpey):

Oops. I meant 1/16. Sorry.

OpenStudy (anonymous):

Isn't it 1/16

OpenStudy (valpey):

Hard to type in LaTeX and keep track of the problem.

OpenStudy (valpey):

How long did it take to get to the post office?

OpenStudy (anonymous):

How do I do that I got 12

OpenStudy (valpey):

You got a Distance of 12 km to the post office. How long did it take to get there?

OpenStudy (anonymous):

Back and forth it took 75 min

OpenStudy (valpey):

Right. So how long did it take to get there?

OpenStudy (valpey):

I'll give you a hint: If the distance was 24 km, it would have taken 1 hour to get there.

OpenStudy (valpey):

Remember: Distance = Speed * Time

OpenStudy (anonymous):

the question is how far does kim live from thhe post office

OpenStudy (valpey):

@Tahmina, I am suggesting that you check your answer. If 12km is correct then we should be able to quickly check that the travel times make sense. This kind of problem can easily end up with impossible solutions if we make a mistake.

OpenStudy (valpey):

She travels 24kmph to the post office, so if the distance is 12 km, it should take how long?